The summers family has 5 childern Patty is 7 years younger than her older brother brother Adam. Tracy and Adam are twins Janelle is Half as old as Patty. Brian who is 13 is 6 years older than Janelle how old is Adam?
Answer by ankor@dixie-net.com(22740)

The summers family has 5 children.
Write an equation for each statement:
:
Patty is 7 years younger than her older brother brother Adam.
P = A - 7
:
Tracy and Adam are twins
T = A
:
Janelle is Half as old as Patty.
J = .5P
:
Brian who is 13 is 6 years older than Janelle
J = 13-6
J = 7
:
From the information find P's age
.5P = 7
P = 14 is P's age
:
how old is Adam?
We know P = A - 7, therefore
A = P + 7
we know that P is 14
A = 14 + 7
A = 21 yrs is Adam's age